### v3

# Who’s important?—

Abstract
As news reporting becomes increasingly global and decentralized online, tracking related events across multiple sources presents significant challenges. Existing news summarization methods typically utilizes Large Language Models and Graphical methods on article-based summaries. However, this is not effective since it only considers the textual content of similarly dated articles to understand the gist of the event. To counteract the lack of analysis on the parties involved, it is essential to come up with a novel framework to gauge the importance of stakeholders and the connection of related events through the relevant entities involved. Therefore, we present SUnSET: Synergistic Understanding of Stakeholder, Events and Time for the task of Timeline Summarization (TLS). We leverage powerful Large Language Models (LLMs) to build SET triplets and introduced the use of stakeholder-based ranking to construct a metric, which can be extended into general situations. Our experimental results outperform all prior baselines and emerged as the new State-of-the-Art, highlighting the impact of stakeholder information within news article.
Who’s important?—
: Synergistic Understanding of Stakeholder,
Events and Time for Timeline Generation
Tiviatis Sim1,2, Yang Kaiwen3, Shen Xin3, Kenji Kawaguchi1 1National University of Singapore, 2A*STAR Institute of High Performance Computing, 3Celia Research Team, Norbert Wiener Research Center, Huawei International tiviatis@u.nus.edu, kaiwen_yang@alumni.brown.edu, sxstar@zju.edu.cn, kenji@comp.nus.edu.sg,
1 Introduction
The abundance of online news media in the advent of the information era has led to a growing volume of daily data production (Statista, ), which poses significant challenges in efficiently identifying and understanding information concisely (Arnold et al., 2023). To promote efficacy and prevent information overloading, automated tasks such as Timeline Summarization (TLS) (Chieu and Lee, 2004; Li et al., 2021; Hu et al., 2024) becomes especially critical in this era (Qorib et al., 2025; Wu et al., 2025). TLS is the generation of a summarized timeline of events, where multiple date-event pairs are sequentially listed to form a narrative for a particular topic. This process generally involves two essential requirements; (1) the events generated must be sequential and succinct (2) only relevant, fundamental and essential events should be included.
Several papers have attempted to improve the brevity of the first requirement and the identification of the second through the use of powerful Large Language Models (LLMs). These methods typically yield good results (Wang et al., 2023; Sojitra et al., 2024a; Hu et al., 2024; Qorib et al., 2025; Wu et al., 2025) due to the emergence of temporal understanding capabilities found within LLMs (Xiong et al., 2024). Nevertheless, the utilization of LLMs in TLS is still relatively underexplored (Hu et al., 2024; Sojitra et al., 2024a) as compared to works on temporal reasoning such as temporal question answering tasks (Piryani et al., 2025).
Although previous works have utilized LLMs to aid temporal understanding within TLS frameworks, they typically focus on article summarization. Moreover, earlier approaches predating LLMs primarily addressed causal and temporal influences. Yet, stakeholders directly involved in events are crucial for identifying links between events and assessing their significance—an aspect largely overlooked in current TLS research. Incorporating stakeholder information can enhance both sequential curation and importance selection by grounding the timeline in real-world relevance.
As such, we propose SUnSET: Synergistic Understanding of Stakeholder, Events and Time, for the task of automated TLS generation (Fig. 1). Unlike prior methods, SUnSET attempts to extract relevant information of various stakeholders mentioned within articles. Moreover, instead of summarizing articles immediately (Qorib et al., 2025; Hu et al., 2024; Sojitra et al., 2024b), SUnSET attempts to extract multiple relevant events mentioned within an article to enable a better representation of occurrences documented by a single news article. The extracted stakeholder and event information will subsequently be combined with its estimated timeframe to generate a Stakeholder-Event-Time triplet (SET). This will be utilized in various clustering steps of the timeline generation process to eventually generate a singular summarized timeline. SUnSET was tested against classic TLS datasets (Tran et al., 2013, 2015) and was able to improve existing baselines to emerge as the new state of the art.
Henceforth, the key contributions from our paper include the following:
1.
A novel framework is introduced to construct and apply SET for TLS.
2.
This is the first work to propose methods for extracting and utilizing stakeholder information within TLS.
3.
Multiple scoring mechanisms are developed, along with constraint-based proofs of stakeholder relevance applicable to general cases.
4.
Empirical evaluations across TLS datasets demonstrate improved performance, validating the benefits of SUnSET in TLS tasks.
2 Related Works
Recent TLS approaches can be broadly categorized into two main types: traditional methods and LLM-augmented methods; although variants of traditional clustering methods are still popular approaches when it comes to TLS task (Gholipour Ghalandari and Ifrim, 2020; Chen et al., 2023), the rise of emergent capabilities in generative LLMs has led to increasing efforts to leverage these tools for TLS tasks, often with diverse objectives. Although LLM applications are more commonly associated with temporal reasoning, their usage in TLS has grown significantly over the past three years. For instance, Wu et al.’s paper highlighted the importance of temporal and causal relationships by incorporating LLMs into self-questioning strategies. On the other hand, works by Hu et al. and Qorib et al. introduced the notion of LLMs functioning as pseudo-oracles, emulating crowdsourced event clustering via pairwise querying during the clustering process.
Clustering methods in TLS are typically implemented in two distinct ways (Gholipour Ghalandari and Ifrim, 2020). The more conventional approach involves grouping key dates to generate graph-based rankings (Steen and Markert, 2019) where subsequent summaries will be formed by selecting the top candidate sentences derived from generated ranked lists of important dates. Alternatively, the other method involves ‘event’ clustering (Ribeiro et al., 2017), where article-level summaries generated once from each article are aggregated based on textual similarity. Increasingly, LLMs are integrated into these processes, especially to enhance summarization. For example, Hu et al. applied LLMs for article-level summarization prior to clustering, followed by a reclustering stage where the LLM validates cluster cohesion through pairwise comparisons of summary nodes. Meanwhile, Wu et al. explored iterative prompting techniques to survey context by progressively self-refining queries with LLMs to retrieve more relevant content. Building upon these efforts, our paper advances TLS clustering with LLM augmentation, where we will expand on stakeholder-based heuristics to assess event relevance.
3 Methodology
We introduce SUnSET: Synergistic Understanding of Stakeholder, Events and Time in this paper for the task of TLS. SUnSET aims to utilize stakeholder information while identifying multiple event occurrences within a single article for creating a graph representation of important events interconnected to each other. The general pipeline can be seen in Figure 2, where the pipeline is split into three main sequences.
3.1 SET Generation
The first step of SUnSET requires the generation of Stakeholder, Event and Time to form a triplet (SET). This is done via utilizing a LLM to extract events mentioned within an article and their estimated date of occurrence. Therefore, every article may contain multiple events: . After extracting the event and its estimated date (), we call the same LLM to extract relevant stakeholders mentioned within the article with a maximum of 5 stakeholders per event: . The term stakeholder used here strictly refers to an entity which is either a person or an organization. This results in a series of SET for a single article: . The relevant prompts used can be found in Appendix A. To minimize disruptions from hallucination, we used heuristic to ensure the generated date and event exists within the article and removed illogical dates. For events with lack of reference dates due to the lack of information within the article, the event uses the date of publication.
Subsequently, Coreference Resolution will be done for all of the extracted stakeholders due to difference in naming such as utilizing a title or a position than a name (E.g. President of the United States) or differences in naming (E.g. POTUS v.s. President of America). A knowledge graph (KG) was built from the wikidata (Vrandečić and Krötzsch, 2014) database. This allows us to capture entities by origin-language forms and transliterations common within non-English dominant regions such as Semitic languages. The specifics of the KG building can be accessed in Appendix B.
3.2 Event Clustering
| Rarity Across Topics | |||
| Reoccurrence | Rare | Common | |
Low |
Normal | Irrelevant | |
High |
Significant | Normal | |
The next stage of the workflow involves the event clustering process. Typically, this involves using content found within similarly dated articles to generate clusters, where each article forms an event node. Instead of using the first five sentences in the body of each article (Gholipour Ghalandari and Ifrim, 2020) or a pre-generated article summary (Hu et al., 2024; Qorib et al., 2025) as the input, SUnSET uses every single event () to generate a more accurate representation of an event’s impact. This clustering process automatically reduces error brought forward by hallucination not caught in the previous step since hallucinations generate small event clusters which would be neglected due to their small event size.
Additionally, unlike the aforementioned clustering methods which uses only date or cosine similarity for cluster creation, we came up with a new metric to gauge the importance of an event. This new metric incorporates stakeholder information; intuitively, the more the stakeholders occurs across all topics , the less important they are. Similarly, if there is frequent reoccurrence of stakeholders within the same topic (), they are more significant (Tab. 1). Therefore, SUnSET clustering introduces the use of Equation 1 to represent the relevance of stakeholders. This score penalizes common stakes across (Eqn. 2) while rewarding within-topic reoccurrence (Eqn. 3).
The penalty score () utilizes the coefficient of variation (Brown, 1998) to grasp the relative dispersion of stakeholder counts with respect to the sample mean across all topics. This value will be max-min normalized, and multiplied by the percentage of occurrence in the current topic. The boundary of and related proof will be in Appendix C. On the other hand, the reward score () uses a dampened hyperbolic tangent score with a factor of to limit the reward value from exploding, where the maximum possible value is capped at . The final relevance score thus multiplies the reward with the penalty score, before using a hyperparameter to tune the significance of the relevance score.
To test the efficacy of the aforementioned penalty score, we introduce another penalty scoring, which is an altered version of the inverse document frequency (IDF) found in BM25 (Robertson and Zaragoza, 2009). Unlike Equation 2, Equation 5 directly uses the absence and presence of stakeholder counts across all documents to estimate the degree to penalize; this value however, fails to capture the actual extent of topic-based counts since it does not consider topic-based absence but document-based absence across the full document set. This may lead to edge cases being misrepresented: cases with common rarity and low reoccurrence (Irrelevant) may have same penalty output as those with both rare and high reoccurrence (Significant) since both cases provides high values. Contrary to using , the coefficient of variation in the original penalty score represents the spread of across all topics, which allows the subsequent penalization of compact counts while identifying low and high frequency topics in high-variation counts through percentage multiplications. The behaviour of the different penalty and reward scores will be further illustrated in Appendix D.
To increase the strictness of relevant events within the same cluster, we have also experimented with incorporating stakeholder Exact Matching (EM), where every 2 nodes requires at least N-matching unique stakeholders to have an edge. Therefore, the clustering process uses the encoded event summary from a General Text Embedding (GTE) Model (Zhang et al., 2024) to obtain query-based cosine similarity scores combined with relevancy scores to generate the top 20 similar events for every node. These events have to satisfy the boolean EM condition to gain a weighted connection () with the query node (Eqn. 4). Due to the use of relevancy scores, our method does not need to incorporate resource-intensive pairwise LLM comparisons of events found in the same cluster (Hu et al., 2024; Qorib et al., 2025).
3.3 Timeline Generation
After obtaining multiple clusters, each cluster undergoes a re-ranking based on both their size and the existing relevance of their events’ stakeholders (Eqn. 7); instead of incorporating relevance scoring to aid in edge connection (Eqn. 4), the relevance score is applied to determine how significant each cluster is relative to both and . Events which are measured as significant will then be passed into TextRank (Mihalcea and Tarau, 2004) to identify important nodes within existing clusters. The final set of nodes will subsequently be used in the final Timeline Generation (TLG).
Henceforth, variations of overlapping events over different articles will not be a pressing concern since the clustering process is able to identify these variations based both the score from similar texts and their scores from repeated stakeholders to create a tighter cluster. In addition, TextRank, which causes nodes which are more different but within the same cluster to be used, would balance the variations in duplicates within their own cluster to better represent the events which are same with differing narrations. The penalty represented in also helps prevent similar but different events from falling into wrong cluster.
4 Experimental Setups
Datasets.
We used two renowned TLS datasets, Timeline17 (T17) (Tran et al., 2013) and Crisis (Tran et al., 2015). T17 contains 19 timelines compiled from varying sources of online news sites, spanning 9 major topics from 2005-2013, each with 1-5 ground truth timelines. On the other hand, Crisis has 22 annotated timelines covering 4 critical crisis events, each containing 4-7 ground truth timelines.
Benchmark.
Models and Deployment.
For experiments, we adopted Qwen2.5-72B-Instruct (Qwen et al., 2025), and GPT-4 omni (OpenAI, 2024). We used GTE-Modernbert-Base for encoding. For the main experimental results, we used the best model, Qwen2.5-72B-Instruct to compare across the different benchmarks. All model deployment were done with VLLM (Kwon et al., 2023) with 8 H200 GPUs. We used batch sizes of 256 for batch encoding, and limit each event summary to a maximum of 4096 characters.
Metrics.
Adhering to previous work done (Hu et al., 2024; Qorib et al., 2025; Wu et al., 2025), we incorporate a part of the Tilse framework (Steen and Markert, 2019) and use three main scores to analyse the performance of the TLS task. We use an Alignment-based ROUGE-1 F1-score (AR-1) to evaluate the semantic distance of unigram overlaps between generated timelines and the provided ground truth. We also used the corresponding bigram overlaps and scored the Alignment-based ROUGE-2 F1-score (AR-2). Lastly, the Date-F1 metric will score the similarity of the dates in the generated timeline compared to the referenced ground truth to understand the quality of major events picked up.
Setups.
Following prior papers (Hu et al., 2024; Qorib et al., 2025), the TLS task can be defined as follows: given a set of temporally labelled news articles that is related to a broad topic (), as well as the expected number of dates and the number of sentences to include in each date, a single timeline summarizing the topic should be generated. This generated timeline will subsequently be compared with the ground truth provided in the datasets. The baselines from the selected benchmarks are run in their publicly available Github repository. We have also published SUnSET publicly on Github for reproduction.
5 Experimental Results
Table 2 showcases SUnSET as compared to the previously SOTA baselines. We selected the strongest model for each method for the comparisons. We also included Llama2-13B in Table 2 since Llama2 was the original model used in LLM-TLS’s paper (Hu et al., 2024) and it occasionally performs better than Qwen2.5-72B.
| Dataset | Method | LLM | AR-1 | AR-2 | Date-F1 |
| Crisis | CHRONOS | Qwen72B | 0.108 | 0.045 | 0.323 |
| LLM-TLS | Llama13B | 0.112 | 0.032 | 0.329 | |
| LLM-TLS | Qwen72B | 0.111 | 0.036 | 0.326 | |
| SUnSET | Qwen72B | 0.129 | 0.047 | 0.389 | |
| SUnSET | GPT-4o | 0.107 | 0.036 | 0.381 | |
| T17 | CHRONOS | Qwen72B | 0.116 | 0.042 | 0.522 |
| LLM-TLS | Llama13B | 0.118 | 0.036 | 0.528 | |
| LLM-TLS | Qwen72B | 0.114 | 0.040 | 0.543 | |
| SUnSET | Qwen72B | 0.136 | 0.044 | 0.576 | |
| SUnSET | GPT-4o | 0.120 | 0.039 | 0.590 |
Our method is able to defeat all prior work in all our metrics used for evaluation. For instance, SUnSET managed to improve Crisis Date-F1 score by , ROUGE-1 by and ROUGE-2 by as compared to the previous SOTA result. The increase in Date-F1 scores over all prior methods indicates the superiority of SUnSET in correctly identifying important dates that were curated in the ground truth, whereas the high ROUGE scores support SUnSET in its ability to retain important details in the distillation process. SUnSET also managed to obtain the best scores in T17, surpassing the previous SOTA. The full set of experimental results are included in Appendix F; examples of timeline output can be accessed in Appendix G. For reference, we attempted using SUnSET with a closed-sourced model (GPT-4omni). GPT-4o performed relatively well and was able to have strong improvements in Date-F1 scores, but lacks in its ROUGE scores as compared to Qwen experiments. To understand where the difference in improvement comes from, more experiments were done on the clustering and timeline generation processes.
5.1 Relevancy in Event Clustering
| Dataset | Method | AR-1 | AR-2 | Date-F1 |
| Crisis | w/o | 0.114 | 0.041 | 0.354 |
| 0.129 | 0.047 | 0.389 | ||
| T17 | w/o | 0.127 | 0.038 | 0.551 |
| 0.136 | 0.044 | 0.576 |
We further investigated the impact of relevance used in the event clustering process by ablating all stakeholder-related information and relevancy scores, retaining only the cosine similarity scores of the events. From the results observed in Table 3, the use of relevance scoring and stakeholder information causes a leap in performance on Date-F1 scores by for Crisis and for T17. This indicates that our method for the identification of stakeholders plays an integral part in ranking event importance.
Moreover, even without implementing , by comparing Table 2’s LLM-TLS score and Table 3’s “w/o ”, it illustrates the significance of the event component (E). It is observed that utilizing LLMs to generate event sets in SET is already more effective than existing methods which typically uses summarization of (Qorib et al., 2025; Hu et al., 2024) or a self-questioning rewriter (Wu et al., 2025). Even without pairwise matching (“w/o ”), using only the cosine similarity of managed to beat the benchmark set by LLM-TLS and CHRONOS (Tab. 2). This reveals that extraction of multiple mentioned events within a single article is vital for TLS tasks; this is intuitive, as news articles often reference related developments that connect to the central event being reported. Significant occurrences may receive limited coverage initially, until a subsequent development—part of an ongoing sequence—draws renewed attention from multiple news outlets. Therefore, since the time component in SET is always a necessary input due to the nature of this task, we have shown that all components in SET are important for the performance of timeline generation.
5.2 Relevancy in Timeline Generation
As stakeholder relevance was utilized in the Timeline Generation process, we compared the difference between the use and absence of to estimate the significance of each cluster before passing them into TextRank . We used the best performing results for both settings (Tab. 4). The inclusion of relevance scoring improved the performance across all metrics for both Crisis and T17; this demonstrates the efficacy of utilizing for cluster ranking, which ultimately improved the selection of important dates while retaining the details of important events.
| Dataset | Method | AR-1 | AR-2 | Date-F1 |
| Crisis | TextRank | 0.117 | 0.040 | 0.368 |
| TextRank+ | 0.129 | 0.047 | 0.389 | |
| T17 | TextRank | 0.128 | 0.041 | 0.559 |
| TextRank+ | 0.136 | 0.044 | 0.576 |
To go a step further, we looked into the influence of the hyperparameter over the entire SUnSET method, where we compared the inclusion and exclusion of during the timeline generation process (Fig. 3). In the full SUnSET process (TextRank+), all three scores (ROUGE-F1, ROUGE-F2 and Date-F1) increases as increases. This behaviour is more latent when is only used in clustering (TextRank only) as a subsequent decrease in performance is observed as increases beyond plateau.
There are several differences in Crisis and T17’s increase as grows. T17, which contains a larger amount of topics but a smaller pool of articles within each topic (larger and smaller ) tends to have a sharp and significant increase with a small at ; although it shows the use of is effective once introduced, the performance may subsequently become stagnant with less significant changes with the increment of . In contrary, Crisis with well established topics containing large pools of articles tend to benefit more as increases after its initial introduction, where small values are not helpful and potentially worsens performance, and larger values provide remarkable growth.
5.3 Penalty PvP: and
| Dataset | Penalty | AR-1 | AR-2 | Date-F1 |
| Crisis | 0.119 | 0.043 | 0.374 | |
| 0.129 | 0.047 | 0.389 | ||
| T17 | 0.137 | 0.043 | 0.576 | |
| 0.136 | 0.044 | 0.576 |
As we introduced a total of two different penalty scores, we tested the difference in the type of penalty used within the relevance score for both datasets. Table 5 shows the difference in using and for SUnSET, where the best result out of all attempted was used. We observed that the impact on T17, which contains smaller datasets and thus lesser stakeholders than Crisis, is smaller; the use of either and causes little differences in T17’s final results, yet it causes significant differences in a larger dataset such as Crisis. This indicates that is more flexible in more cases, especially when there exist a larger event space where more Significant and Irrelevant stakeholders (Tab. 1) can be identified.
5.4 Stakeholder Exact Matching
| Dataset | EM | AR-1 | AR-2 | Date-F1 | |
| Crisis | 0.114 | 0.041 | 0.354 | ||
| 0.119 | 0.044 | 0.364 | |||
| 0.114 | 0.041 | 0.358 | |||
| 1.0 | 0.129 | 0.047 | 0.389 | ||
| 0.9 | 0.127 | 0.046 | 0.386 | ||
| 0.6 | 0.121 | 0.042 | 0.379 | ||
| T17 | 0.127 | 0.038 | 0.553 | ||
| 0.128 | 0.040 | 0.553 | |||
| 0.127 | 0.040 | 0.558 | |||
| 1.0 | 0.134 | 0.043 | 0.574 | ||
| 0.9 | 0.136 | 0.044 | 0.576 | ||
| 1.0 | 0.134 | 0.044 | 0.573 |
Our initial experiments used only EM with SET without consideration of relevance scores. In Table 6, when is set as 0, there is no relevance scoring used, and we observe that the inclusion of EM is able to capture meaningful stakeholder information to a proportional degree. Typically, is sufficient to result in an improvement, and further matching counts may instead deteriorate the performance of models. This may be due to insufficient stakeholders with high counts, where the median of repetitive stakeholders typically lies around 3 across all datasets (Appn. E). Though EM performance does not match up to the addition of , this is sufficient to signify the importance of stakeholder information in the TLS.
Furthermore, the improvement of over EM illustrates the strength of our formulated equation in stakeholder ranking. While is used, the use of EM causes limited improvements as in itself is able to capture and rank the entities involved in . It is also observed that high values typically performs better across differing EM settings, which further emphasizes the importance of .
5.5 Time Comparison
| Dataset | Method | GPU Time |
| Crisis | LLM-TLS | 7 hours 12 min |
| CHRONOS | 24 min | |
| SUnSET | 7 min 14 sec | |
| T17 | LLM-TLS | 2 hours 12 min |
| CHRONOS | 1 hour 9 min | |
| SUnSET | 1 min 41 sec |
Lastly, we measured the time required for running the experiments, and realised the high efficiency for using SUnSET as compared to previous SOTA methods. The benefits mainly comes from the lack of additional LLM prompting beyond SET creation, where CHRONOS utilizes questioning rounds and LLM-TLS prompts once for every pairwise comparison required. Therefore, SUnSET is able to bypass such time-consuming prompt-questioning methods due to the high efficacy of utilizing stakeholder information for distilling events into a compact representation.
6 Conclusions
We propose SUnSET, a novel approach that leverages SET triplets and stakeholder relevance to generate milestone events for TLS problems. By incorporating stakeholder-relevancy heuristics, SUnSET effectively addresses the challenges of distilling multiple event sets from complex articles. Our framework systematically tackles these issues in three critical steps: 1) SET generation expands the pool of candidate events per article; 2) SUnSET’s efficient event clustering eliminates the need for time-intensive reclustering, even with larger candidate sets; 3) Stakeholder information encoding via enhances both clustering and timeline generation by yielding a more representative set of linked events. Through extensive experiments and ablation studies on the T17 and Crisis datasets, SUnSET consistently delivers state-of-the-art results, validating its effectiveness in both performance and interpretability. These findings underscore the potential of stakeholder-aware TLS systems to advance narrative coherence, relevance, and human-aligned summarization across evolving news domains, while also establishing stakeholder scoring as a scalable framework for prioritizing information in other contexts where audience-centric relevance and societal impact are key — such as search, recommendation, and policy communication.
Limitations
Our study advances TLS through the integration of SET and stakeholder ranking mechanisms, though certain limitations remain. Within our domain of control, we applied feasible refinement techniques such as prompt engineering and performed rigorous reviews to minimize inconsistencies. However, as with prior work, we were unable to fully mitigate hallucinations or factual inaccuracies that may arise during event extraction when utilizing LLMs due to their inherent generation behavior. Finally, in line with previous studies, we evaluated SUnSET using publicly available research datasets, without extending analysis to real-time news data with more dynamic, evolving topics. We identify this as a promising direction for future exploration.
Acknowledgments
We would like to thank Austen Jeremy Sugiarto and Leong Chee Kai in aiding the initial stages of the experiments. We are also grateful to Dr. Basura Fernando for his help in reading through the paper. We deeply appreciate any reviewers for their helpful feedback.
Appendix A Prompts used for LLMs
Appendix B Building Knowledge Graph from Wikidata111Main API: https://www.wikidata.org/wiki/Wikidata:REST_API, NER module: https://huggingface.co/spacy/en_core_web_trf, Request module: https://requests.readthedocs.io/en/latest/
Appendix C Boundary for Penalty Score
Appendix D Behaviour of Reward and Penalty Scores
We illustrate the behaviour of both Penalty and Reward in this section. From Figure 4, Graph plots (Eqn. 3) as the count of stakeholder increases. It can be observed that when the count of the stakeholders () reach approximately 21, the value of saturates regardless of any more addition to counts for the same stakeholder. Subsequently, Graph plots the equation of unnormalized (Eqn. 2) of different scenarios (Tab. 8).
[…]
| 0 | A: 2, B: 3, C:5 | Close Distribution, All Low Counts |
| 1 | A: 90, B: 85, C:65 | Close Distribution, All High Counts |
| 2 | A: 5, B: 5, C:5 | Uniform Distribution, All Low Counts |
| 3 | A: 16, B: 16, C:16 | Uniform Distribution, All High Counts |
| 4 | A: 15, B: 4, C:54 | Single-Peak Distribution, Peak High Count |
| 5 | A: 1, B: 8, C:2 | Single-Peak Distribution, Peak Low Counts |
| 6 | A: 6, B: 7, C:1 | Double-Peaks Distribution, Peaks Low Counts |
| 7 | A: 21, B: 19, C:3 | Double-Peaks Distribution, Peaks High Counts |
| 8 | A: 3, B: 0, C:0 | Single-Peak Distribution, Peak Low Count, Rest are 0 |
| 9 | A: 19, B: 0, C:0 | Single-Peak Distribution, Peak High Count, Rest are 0 |
| 10 | A: 6, B: 3, C:3 | Single-Peak Distribution, Peak Low Count, Rest are half |
| 11 | A: 26, B: 13, C:13 | Single-Peak Distribution, Peak High Count, Rest are half |
Graph plots the value of without consideration of (by setting ) to understand the impact of multiplying and . Cases 0, 1, 2, 3, 10 and 11 illustrates when stakeholder counts are well spread across the topics, their values are low irregardless of high or low counts, which is the desired behaviour since stakeholder with such behaviours fall under either ‘Normal’ or ‘Irrelevant’ (Tab. 1). Alternatively, ‘Significant’ cases with strong single-peak distribution dominating over the others (Indices 4, 5, 8, 9) show significantly higher values, where completely unique stakeholders tend to reach maximum possible . It is noteworthy to add that the difference between the final score between cases 8 and 9 is important since the scores are able to dampen case 8’s score due to its low count value which was set to the common median (3) for illustration. This also occurs between cases 4 and 5. Lastly, those with double-peak distribution (Indices 6 and 7) have a mid-low value, where the duo peaks’ are significantly higher than that of the topic with the low-value count, yet not as high as what would occur in a single-peak. The duo-peaks were penalized with the number of similarly ‘Significant’ topics within themselves. This leads to a lower overall score.
Similarly, we will examine the behaviour of . Graph 1 of Figure 5 also shows for reference. Graph 2 plots unnormalized value to show how it changes as increases (Eqn. 5). It should be noted that indices 1, 3, 4, 7, 9 and 11 (Tab. 8) cannot be differentiated between topics (i.e. A, B and C) since there is no differentiation between topics while using which makes it inferior to . Despite so, it can be observed from Graph 2 that as more articles uses the same stakeholder, becomes smaller, mirroring the way inverse document frequencies work.
Graph 3 further documents the interaction of and , where every line shows how the differing counts of stakeholders across topic () interact when multiplies which rewards within a topic (). It is observed that the relevance score of ‘Irrelevant’ cases (Tab. 1) will still be highly rewarded beyond 1 once goes beyond 5 regardless of , for instance having 5 out of 100 of any stakeholder allows a of around 1, which is not ideal despite its proportionally low count value. This flaw from shows that proportion and distribution should be essential, which is accounted for within .
Appendix E Repeating Stakeholder Counts
We recorded the counts for each stakeholder across all topics in our dataset. There are occasional outliers where the entity involved may be mentioned more than a hundred times and may even reach the thousands, but typically, a count of ten or less is normal across all stakeholders existing in all topics. From Figures 7 and 7, we can observe that the individual median of each topic regardless of the dataset is around three. This means that within their own topics, most stakeholders are mentioned at least in three different articles.
Appendix F Full Set of Experiment Results
| TextRank+ | TextRank | ||||||
| beta | EM | AR-1 | AR-2 | Date-F1 | AR-1 | AR-2 | Date-F1 |
| 0 | 0 | 0.114358 | 0.041683 | 0.354251 | 0.114358 | 0.041683 | 0.354251 |
| 0.1 | 0 | 0.113652 | 0.039741 | 0.35494 | 0.112811 | 0.039234 | 0.353606 |
| 0.2 | 0 | 0.116541 | 0.041465 | 0.360721 | 0.113745 | 0.039872 | 0.354645 |
| 0.3 | 0 | 0.118841 | 0.043264 | 0.367969 | 0.11459 | 0.040077 | 0.358176 |
| 0.4 | 0 | 0.119222 | 0.043375 | 0.367698 | 0.113476 | 0.039522 | 0.357489 |
| 0.5 | 0 | 0.122711 | 0.04451 | 0.377522 | 0.116135 | 0.039776 | 0.366623 |
| 0.6 | 0 | 0.12393 | 0.04462 | 0.381933 | 0.116748 | 0.040169 | 0.368 |
| 0.7 | 0 | 0.125512 | 0.0456 | 0.384202 | 0.116847 | 0.040308 | 0.368 |
| 0.8 | 0 | 0.127489 | 0.047185 | 0.383934 | 0.116723 | 0.040229 | 0.364686 |
| 0.9 | 0 | 0.12894 | 0.047628 | 0.387541 | 0.116548 | 0.040077 | 0.364686 |
| 1 | 0 | 0.129269 | 0.047488 | 0.388972 | 0.116641 | 0.04008 | 0.364686 |
| 0 | 1 | 0.119383 | 0.044027 | 0.364554 | 0.119383 | 0.044027 | 0.364554 |
| 0.1 | 1 | 0.11499 | 0.040711 | 0.358591 | 0.114242 | 0.040103 | 0.357267 |
| 0.2 | 1 | 0.117851 | 0.04203 | 0.363621 | 0.115888 | 0.041408 | 0.361055 |
| 0.3 | 1 | 0.119364 | 0.042948 | 0.368201 | 0.115872 | 0.041243 | 0.361055 |
| 0.4 | 1 | 0.119155 | 0.042993 | 0.37158 | 0.115795 | 0.04119 | 0.360607 |
| 0.5 | 1 | 0.120324 | 0.043291 | 0.374345 | 0.116859 | 0.040401 | 0.366181 |
| 0.6 | 1 | 0.123812 | 0.044156 | 0.385318 | 0.116893 | 0.040489 | 0.366181 |
| 0.7 | 1 | 0.124445 | 0.044881 | 0.383322 | 0.117104 | 0.040751 | 0.366181 |
| 0.8 | 1 | 0.125976 | 0.04591 | 0.383054 | 0.117077 | 0.040633 | 0.366181 |
| 0.9 | 1 | 0.127911 | 0.046209 | 0.386313 | 0.116135 | 0.039776 | 0.366623 |
| 1 | 1 | 0.124948 | 0.045573 | 0.374559 | 0.114044 | 0.039693 | 0.366867 |
| 0 | 2 | 0.114743 | 0.041232 | 0.358151 | 0.114743 | 0.041232 | 0.358151 |
| 0.1 | 2 | 0.114549 | 0.039838 | 0.365324 | 0.113659 | 0.039767 | 0.361014 |
| 0.2 | 2 | 0.115156 | 0.039829 | 0.367784 | 0.113425 | 0.039519 | 0.361014 |
| 0.3 | 2 | 0.118806 | 0.041571 | 0.371988 | 0.114599 | 0.040124 | 0.361038 |
| 0.4 | 2 | 0.120898 | 0.04296 | 0.37482 | 0.115251 | 0.040805 | 0.363499 |
| 0.5 | 2 | 0.121954 | 0.043179 | 0.377068 | 0.115556 | 0.040812 | 0.363499 |
| 0.6 | 2 | 0.121427 | 0.04293 | 0.379509 | 0.116569 | 0.040968 | 0.365772 |
| 0.7 | 2 | 0.120243 | 0.042156 | 0.374532 | 0.1158 | 0.040556 | 0.364149 |
| 0.8 | 2 | 0.120882 | 0.042673 | 0.375772 | 0.115812 | 0.040559 | 0.364149 |
| 0.9 | 2 | 0.122365 | 0.043452 | 0.375772 | 0.115122 | 0.04015 | 0.365345 |
| 1 | 2 | 0.12309 | 0.043711 | 0.375196 | 0.115122 | 0.040148 | 0.365345 |
| TextRank+ | TextRank | ||||||
| beta | EM | AR-1 | AR-2 | Date-F1 | AR-1 | AR-2 | Date-F1 |
| 0 | 0 | 0.127151 | 0.038522 | 0.551891 | 0.127151 | 0.038522 | 0.551891 |
| 0.1 | 0 | 0.133391 | 0.042983 | 0.567552 | 0.131371 | 0.042254 | 0.554819 |
| 0.2 | 0 | 0.132674 | 0.042788 | 0.566986 | 0.128913 | 0.040551 | 0.550866 |
| 0.3 | 0 | 0.13343 | 0.04304 | 0.569759 | 0.128958 | 0.040486 | 0.556219 |
| 0.4 | 0 | 0.13434 | 0.043684 | 0.571095 | 0.128798 | 0.040446 | 0.556219 |
| 0.5 | 0 | 0.136108 | 0.044275 | 0.571878 | 0.129042 | 0.040489 | 0.556219 |
| 0.6 | 0 | 0.135763 | 0.044037 | 0.571878 | 0.12782 | 0.03994 | 0.55434 |
| 0.7 | 0 | 0.13604 | 0.044052 | 0.569624 | 0.128425 | 0.040278 | 0.554834 |
| 0.8 | 0 | 0.13582 | 0.043954 | 0.571009 | 0.128086 | 0.040158 | 0.554834 |
| 0.9 | 0 | 0.134498 | 0.043165 | 0.574814 | 0.127756 | 0.03997 | 0.554834 |
| 1 | 0 | 0.134713 | 0.043138 | 0.574814 | 0.127738 | 0.039948 | 0.554834 |
| 0 | 1 | 0.128352 | 0.040235 | 0.553264 | 0.127691 | 0.039799 | 0.551384 |
| 0.1 | 1 | 0.134273 | 0.043627 | 0.56971 | 0.131719 | 0.042692 | 0.553908 |
| 0.2 | 1 | 0.133477 | 0.043096 | 0.571774 | 0.129177 | 0.041106 | 0.550931 |
| 0.3 | 1 | 0.134157 | 0.043373 | 0.571774 | 0.129145 | 0.041125 | 0.55464 |
| 0.4 | 1 | 0.134424 | 0.043731 | 0.571774 | 0.128781 | 0.040995 | 0.55464 |
| 0.5 | 1 | 0.136009 | 0.044221 | 0.572557 | 0.128841 | 0.040975 | 0.55464 |
| 0.6 | 1 | 0.136132 | 0.044262 | 0.572557 | 0.127743 | 0.040477 | 0.549568 |
| 0.7 | 1 | 0.136607 | 0.044352 | 0.571172 | 0.128484 | 0.041047 | 0.550062 |
| 0.8 | 1 | 0.136573 | 0.044345 | 0.573073 | 0.128222 | 0.040971 | 0.550062 |
| 0.9 | 1 | 0.136085 | 0.043953 | 0.576362 | 0.128222 | 0.040947 | 0.550062 |
| 1 | 1 | 0.136176 | 0.043886 | 0.576362 | 0.128222 | 0.040947 | 0.550062 |
| 0 | 2 | 0.127571 | 0.04065 | 0.558906 | 0.127571 | 0.04065 | 0.558906 |
| 0.1 | 2 | 0.134279 | 0.044326 | 0.567725 | 0.127598 | 0.041011 | 0.548946 |
| 0.2 | 2 | 0.132227 | 0.04378 | 0.565846 | 0.126087 | 0.039899 | 0.548946 |
| 0.3 | 2 | 0.13283 | 0.043909 | 0.57041 | 0.126327 | 0.03999 | 0.55172 |
| 0.4 | 2 | 0.134523 | 0.04468 | 0.57041 | 0.127085 | 0.040477 | 0.55172 |
| 0.5 | 2 | 0.133141 | 0.043959 | 0.57041 | 0.127483 | 0.040714 | 0.55172 |
| 0.6 | 2 | 0.133539 | 0.044212 | 0.57041 | 0.127142 | 0.040493 | 0.55172 |
| 0.7 | 2 | 0.133565 | 0.043979 | 0.57041 | 0.127077 | 0.040308 | 0.55172 |
| 0.8 | 2 | 0.133548 | 0.043975 | 0.571507 | 0.126337 | 0.040098 | 0.55172 |
| 0.9 | 2 | 0.133639 | 0.04401 | 0.571507 | 0.124933 | 0.038988 | 0.54984 |
| 1 | 2 | 0.13477 | 0.044337 | 0.573387 | 0.124747 | 0.039042 | 0.54984 |
| TextRank+ | TextRank | ||||||
| beta | EM | AR-1 | AR-2 | Date-F1 | AR-1 | AR-2 | Date-F1 |
| 0 | 0 | 0.114358 | 0.041683 | 0.354251 | 0.114358 | 0.041683 | 0.354251 |
| 0.1 | 0 | 0.119213 | 0.042284 | 0.364802 | 0.11779 | 0.040881 | 0.363209 |
| 0.2 | 0 | 0.11937 | 0.042153 | 0.365167 | 0.117337 | 0.040573 | 0.363209 |
| 0.3 | 0 | 0.119812 | 0.042737 | 0.369653 | 0.117364 | 0.040541 | 0.363942 |
| 0.4 | 0 | 0.118798 | 0.042092 | 0.370735 | 0.118033 | 0.041105 | 0.367189 |
| 0.5 | 0 | 0.119503 | 0.043181 | 0.375763 | 0.118011 | 0.041095 | 0.367189 |
| 0.6 | 0 | 0.118468 | 0.042296 | 0.37138 | 0.118011 | 0.041095 | 0.367189 |
| 0.7 | 0 | 0.118511 | 0.042324 | 0.370646 | 0.118011 | 0.041095 | 0.367189 |
| 0.8 | 0 | 0.118789 | 0.042531 | 0.369889 | 0.118011 | 0.041095 | 0.367189 |
| 0.9 | 0 | 0.119232 | 0.043089 | 0.373136 | 0.118011 | 0.041095 | 0.367189 |
| 1 | 0 | 0.118836 | 0.042856 | 0.373136 | 0.118011 | 0.041095 | 0.367189 |
| 0 | 1 | 0.119383 | 0.044027 | 0.364554 | 0.119383 | 0.044027 | 0.364554 |
| 0.1 | 1 | 0.116948 | 0.041352 | 0.368236 | 0.11742 | 0.041117 | 0.3668 |
| 0.2 | 1 | 0.114585 | 0.040833 | 0.357094 | 0.117598 | 0.041219 | 0.3668 |
| 0.3 | 1 | 0.114909 | 0.041108 | 0.362224 | 0.11765 | 0.041245 | 0.367534 |
| 0.4 | 1 | 0.116967 | 0.041394 | 0.371358 | 0.117646 | 0.041252 | 0.367534 |
| 0.5 | 1 | 0.117696 | 0.042491 | 0.373769 | 0.117646 | 0.041252 | 0.367534 |
| 0.6 | 1 | 0.117333 | 0.042068 | 0.37128 | 0.117646 | 0.041252 | 0.367534 |
| 0.7 | 1 | 0.119144 | 0.043011 | 0.374527 | 0.117646 | 0.041252 | 0.367534 |
| 0.8 | 1 | 0.118695 | 0.042807 | 0.373769 | 0.117646 | 0.041252 | 0.367534 |
| 0.9 | 1 | 0.118588 | 0.042762 | 0.373769 | 0.117646 | 0.041252 | 0.367534 |
| 1 | 1 | 0.117075 | 0.04204 | 0.372349 | 0.117646 | 0.041252 | 0.367534 |
| 0 | 2 | 0.114743 | 0.041232 | 0.358151 | 0.114743 | 0.041232 | 0.358151 |
| 0.1 | 2 | 0.113404 | 0.039178 | 0.366255 | 0.112941 | 0.039047 | 0.363297 |
| 0.2 | 2 | 0.112839 | 0.039618 | 0.360665 | 0.112568 | 0.03918 | 0.361919 |
| 0.3 | 2 | 0.112525 | 0.038941 | 0.361231 | 0.113128 | 0.039125 | 0.364192 |
| 0.4 | 2 | 0.112189 | 0.038723 | 0.362651 | 0.113128 | 0.039093 | 0.364192 |
| 0.5 | 2 | 0.113626 | 0.039906 | 0.364744 | 0.113128 | 0.039093 | 0.364192 |
| 0.6 | 2 | 0.112348 | 0.039172 | 0.363323 | 0.112722 | 0.03876 | 0.364192 |
| 0.7 | 2 | 0.112828 | 0.03947 | 0.364744 | 0.11269 | 0.038761 | 0.364192 |
| 0.8 | 2 | 0.112622 | 0.039416 | 0.364281 | 0.11269 | 0.038761 | 0.364192 |
| 0.9 | 2 | 0.112912 | 0.039573 | 0.364281 | 0.11269 | 0.038761 | 0.364192 |
| 1 | 2 | 0.113113 | 0.039672 | 0.363477 | 0.11269 | 0.038761 | 0.364192 |
| TextRank+ | TextRank | ||||||
| beta | EM | AR-1 | AR-2 | Date-F1 | AR-1 | AR-2 | Date-F1 |
| 0 | 0 | 0.127151 | 0.038522 | 0.551891 | 0.127151 | 0.038522 | 0.551891 |
| 0.1 | 0 | 0.133361 | 0.042997 | 0.571505 | 0.128408 | 0.040082 | 0.55271 |
| 0.2 | 0 | 0.133766 | 0.042873 | 0.575374 | 0.128416 | 0.039956 | 0.554628 |
| 0.3 | 0 | 0.134439 | 0.042803 | 0.575513 | 0.128083 | 0.039733 | 0.556272 |
| 0.4 | 0 | 0.135246 | 0.043226 | 0.575513 | 0.128149 | 0.039836 | 0.556272 |
| 0.5 | 0 | 0.135214 | 0.043271 | 0.574128 | 0.128149 | 0.039836 | 0.556272 |
| 0.6 | 0 | 0.135432 | 0.043603 | 0.573612 | 0.128149 | 0.039836 | 0.556272 |
| 0.7 | 0 | 0.135341 | 0.043572 | 0.575207 | 0.128149 | 0.039836 | 0.556272 |
| 0.8 | 0 | 0.134963 | 0.043413 | 0.575658 | 0.12811 | 0.039807 | 0.556272 |
| 0.9 | 0 | 0.13518 | 0.043546 | 0.574529 | 0.12811 | 0.039807 | 0.556272 |
| 1 | 0 | 0.133889 | 0.042507 | 0.57265 | 0.12811 | 0.039807 | 0.556272 |
| 0 | 1 | 0.128352 | 0.040235 | 0.553264 | 0.127691 | 0.039799 | 0.551384 |
| 0.1 | 1 | 0.137057 | 0.044288 | 0.574131 | 0.129329 | 0.040677 | 0.545189 |
| 0.2 | 1 | 0.136771 | 0.043724 | 0.576053 | 0.129147 | 0.040417 | 0.548139 |
| 0.3 | 1 | 0.13739 | 0.043633 | 0.576029 | 0.128679 | 0.040163 | 0.548139 |
| 0.4 | 1 | 0.137267 | 0.043609 | 0.576029 | 0.129101 | 0.040205 | 0.549784 |
| 0.5 | 1 | 0.137116 | 0.043592 | 0.574128 | 0.129101 | 0.040205 | 0.549784 |
| 0.6 | 1 | 0.137352 | 0.043926 | 0.573612 | 0.129101 | 0.040205 | 0.549784 |
| 0.7 | 1 | 0.136805 | 0.043857 | 0.573612 | 0.129101 | 0.040205 | 0.549784 |
| 0.8 | 1 | 0.136675 | 0.043737 | 0.572515 | 0.129101 | 0.040205 | 0.549784 |
| 0.9 | 1 | 0.136709 | 0.043784 | 0.571386 | 0.129101 | 0.040205 | 0.549784 |
| 1 | 1 | 0.135355 | 0.04276 | 0.569507 | 0.129101 | 0.040205 | 0.549784 |
| 0 | 2 | 0.127571 | 0.04065 | 0.558906 | 0.127571 | 0.04065 | 0.558906 |
| 0.1 | 2 | 0.134797 | 0.043389 | 0.574589 | 0.126857 | 0.040582 | 0.551637 |
| 0.2 | 2 | 0.134322 | 0.043568 | 0.57271 | 0.125526 | 0.039409 | 0.553047 |
| 0.3 | 2 | 0.134659 | 0.043783 | 0.573226 | 0.125708 | 0.039558 | 0.553047 |
| 0.4 | 2 | 0.134008 | 0.042938 | 0.57271 | 0.125539 | 0.039507 | 0.552531 |
| 0.5 | 2 | 0.13373 | 0.043024 | 0.57271 | 0.12594 | 0.039747 | 0.552531 |
| 0.6 | 2 | 0.13373 | 0.043039 | 0.57271 | 0.12594 | 0.039747 | 0.552531 |
| 0.7 | 2 | 0.134293 | 0.043185 | 0.57271 | 0.12594 | 0.039747 | 0.552531 |
| 0.8 | 2 | 0.134293 | 0.043185 | 0.57271 | 0.12594 | 0.039747 | 0.552531 |
| 0.9 | 2 | 0.134407 | 0.043212 | 0.57271 | 0.12594 | 0.039747 | 0.552531 |
| 1 | 2 | 0.134075 | 0.043106 | 0.57271 | 0.12594 | 0.039747 | 0.552531 |
Appendix G Example Timelines Generated from Experiments
-2.3cm-2.3cm